# Multi-Turn Interactions in Large Language Models

### v1

@@ 1 Introduction
We model health coaching as a user-conditioned partially observable control problem with typed rewards. The per-turn utility is a user-specific composition defined in Eq. 1, where $z_{t}$ summarizes dialogue context, recent outcomes, and user characteristics (including health literacy and self-efficacy), and the tool component derives from success/failure outcomes. To address cold-start and sparse feedback, we add a small, early-turn user-model information-gain bonus (curiosity) that encourages actions which reduce uncertainty over latent interaction preferences (e.g., tolerance for explanation, desire for analysis). We define this bonus formally in Eq. 2 and apply it only in the first few turns [1, 2, 3, 4].
First, OPE with self-normalized importance sampling and doubly-robust estimators [e.g., 5, 6, 7] lets us compare counterfactual policies on real logs without additional user trials. We find that uniform, heavy tool usage can maximize average reward on logs but causes harm for specific archetypes (e.g., low-literacy/high-efficacy), strengthening the case for archetype-aware reward weighting. Second, in simulation with hidden archetypes, brief early information-gain phase reliably shortens trait-identification time and improves task success and reliability (pass@3), which supports a "probe-then-personalize" strategy consistent with our deployment findings.
@@ 2 Related Work
LLM-based agents have advanced health applications in coaching and diagnostics. PH-LLM [8] fine-tunes models on wearable data for personalized sleep and fitness insights, while PHIA [9] uses iterative reasoning with tools for health data interpretation across multiple turns. Recent work explores behavior-theoretic scaffolding [10] and proactive strategies [11] for long-term coaching. However, multi-turn deployment remains challenging: [12] report substantial degradation over extended conversations, and tool-enabled settings can be brittle. Notably, [13] show that while tools and natural-language feedback generally help performance, multi-turn pipelines remain fragile and certain training choices can hurt longer conversations. This suggests that "more tools is always better" is too simplistic.
@@ 3.1 System and Data
We built and deployed a web-based LLM health system (Qwen3-235B-A22B via API providers [21]) for multi-turn health coaching with real users and wearable data. The system processes user-uploaded health data (e.g., Apple Health) through: (1) a data pipeline that preprocesses historical data into daily features (sleep, HRV, VO2max, activity) and generates ML predictions for stress, soreness, and injury risk (rolling-origin cross-validation; $R^{2}{=}\,0.50,\,0.28,\,0.40$ respectively on 1,280 day-samples, N=28N{=}28) [22]; (2) agent tools including Code Executor, Web Searcher, and Email (reminders and events).
@@ 3.5 Offline Policy Evaluation (OPE)
For each decision head we fit a probabilistic behavior model $\hat{\pi}_{b}^{h}(a^{h}\mid x_{t})$ on features $x_{t}=(z_{t},u_{i})$ (turn features and user attributes) to approximate logging propensities. We assess calibration via Expected Calibration Error (ECE) per head. For primary reporting we use self-normalized importance sampling (SNIPS)[25] for the objective component and augmented inverse propensity weighting (AIPW)[26] for satisfaction.
@@
We model the logging policy as head-factorized over $\{\textsc{Tool},\textsc{Style}\}$ and fit cross-fitted, calibrated classifiers for each head to obtain per-head propensities $\hat{\pi}_{b}^{h}(a_{t}^{h}\mid x_{t})$. Assuming head-wise factorization conditional on $x_{t}$, the joint logging propensity is $\hat{\pi}_{b}(a_{t}\mid x_{t})=\prod_{h}\hat{\pi}_{b}^{h}(a_{t}^{h}\mid x_{t})$, and the target policy factorizes analogously. For SNIPS and selection-aware SNIPS we use the joint importance ratio
@@
with ratio clipping at $c=50$ and session-level bootstrap CIs.
@@
For interpretability we report an unweighted objective $R_{\text{obj}}=R_{\text{tool}}+R_{\text{eng}}$ via SNIPS, alongside a Total that applies fixed, pre-specified literacy-aware weights to form $R_{\text{total}}=\alpha_{w}R_{\text{user}}+\beta_{w}R_{\text{tool}}+\gamma_{w}R_{\text{eng}}$. Here $(\alpha_{w},\beta_{w},\gamma_{w})$ is a deterministic function of literacy strata (low-literacy: $(0.6,0.2,0.2)$; high-literacy: $(0.3,0.5,0.2)$). Because $R_{\text{obj}}$ and Total use different normalizations and estimators (SNIPS for objective components; selection-aware SNIPS/AIPW for satisfaction), columns do not sum algebraically.
@@ 4.4 Counterfactual Policy Analysis
Table 5: Offline policy evaluation on deployment logs (session-bootstrap 95% CIs). $R_{\text{obj}}$ reports a SNIPS estimate of the unweighted objective $R_{\text{tool}}{+}R_{\text{eng}}$; $R_{\text{user}}$ reports an AIPW estimate of satisfaction (z-score). Total == SNIPS of the weighted objective $(\beta_{w}R_{\text{tool}}{+}\gamma_{w}R_{\text{eng}})$ ++ selection-aware SNIPS/AIPW of the $\alpha_{w}$-weighted satisfaction. Weights are literacy-based: low-literacy $(\alpha_{w},\beta_{w},\gamma_{w})=(0.6,0.2,0.2)$; high-literacy $(0.3,0.5,0.2)$. Because the two SNIPS components use different normalizations, columns do not sum algebraically to Total.
Policy $R_{\text{obj}}$ (SNIPS, unweighted) $R_{\text{user}}$ (AIPW) Total Reward

### v2

@@ 1 Introduction
We model health coaching as a user-conditioned partially observable Markov decision process (POMDP) with typed rewards. The per-turn utility is a user-specific composition defined in Eq. 1, where $z_{t}$ summarizes dialogue context, recent outcomes, and user characteristics (including health literacy and self-efficacy), and the tool component derives from success/failure outcomes. To address cold-start and sparse feedback, we add a small, early-turn user-model information-gain bonus (curiosity) that encourages actions which reduce uncertainty over latent interaction preferences (e.g., tolerance for explanation, desire for analysis) [1, 2, 3, 4]. We define this bonus formally in Eq. 2 and apply it only in the first few turns.
First, OPE with self‑normalized IPS (SNIPS) and AIPW (doubly robust) [5, 6, 7] lets us compare counterfactual policies on real logs without additional user trials. We find that uniform, heavy tool usage can maximize average reward on logs but causes harm for specific archetypes (e.g., low-literacy/high-efficacy), strengthening the case for archetype-aware reward weighting. Second, in simulation with hidden archetypes, brief early information-gain phase reliably shortens trait-identification time and improves task success and reliability (pass@3), which supports a "probe-then-personalize" strategy consistent with our deployment findings.
@@ 2 Related Work
LLM-based agents have advanced health applications in coaching and diagnostics. PH-LLM [8] fine-tunes models on wearable data for personalized sleep and fitness insights, while PHIA [9] uses iterative reasoning with tools for health data interpretation across multiple turns. Recent work explores behavior-theoretic scaffolding [10] and proactive strategies [11] for long-term coaching. However, multi-turn deployment remains challenging: [12] report substantial degradation over extended conversations, and tool-enabled settings can be brittle. Notably, [13] show that while tools and natural-language feedback generally help performance, multi-turn pipelines remain fragile and certain training choices can hurt longer conversations.
@@ 3.1 System and Data
We built and deployed a web-based LLM health system (Qwen3-235B-A22B via API providers [21]) for multi-turn health coaching with real users and wearable data. The system processes user-uploaded health data (i.e., Apple Health app export) through: (1) a data pipeline that preprocesses historical data into daily features (sleep, HRV, VO2max, activity) and generates ML predictions for stress, soreness, and injury risk (rolling-origin cross-validation; $R^{2}{=}\,0.50,\,0.28,\,0.40$ respectively on 1,280 day-samples, N=28N{=}28) [22]; (2) agent tools including Code Executor, Web Searcher, and Email (reminders and events).
@@ 3.5 Offline Policy Evaluation (OPE)
For each decision head we fit a probabilistic behavior model $\hat{\pi}_{b}^{h}(a^{h}\mid x_{t})$ on features $x_{t}=(z_{t},u_{i})$ (turn features and user attributes) to approximate logging propensities. We assess calibration via Expected Calibration Error (ECE) per head. For primary reporting we use self-normalized importance sampling (SNIPS)[25] for the objective component and augmented inverse propensity weighting (AIPW) for satisfaction.
@@
We model the logging policy as head-factorized over $\{\textsc{Tool},\textsc{Style}\}$ and fit cross-fitted, calibrated classifiers for each head to obtain per-head propensities $\hat{\pi}_{b}^{h}(a_{t}^{h}\mid x_{t})$. Assuming head-wise factorization conditional on $x_{t}$, the joint logging propensity is $\hat{\pi}_{b}(a_{t}\mid x_{t})=\prod_{h}\hat{\pi}_{b}^{h}(a_{t}^{h}\mid x_{t})$, and the target policy factorizes analogously. For both the SNIPS estimator and the augmentation term in the AIPW estimator, we use the same joint importance ratio
@@
with ratio clipping at c=50c{=}50 and session-level bootstrap confidence intervals.
@@
For interpretability we report an unweighted objective $R_{\text{obj}}=R_{\text{tool}}+R_{\text{eng}}$ with SNIPS, alongside a $R_{\text{total}}$ that applies fixed, pre-specified literacy-aware weights to form $R_{\text{total}}=\alpha_{w}R_{\text{user}}+\beta_{w}R_{\text{tool}}+\gamma_{w}R_{\text{eng}}$. Here $(\alpha_{w},\beta_{w},\gamma_{w})$ is a deterministic function of literacy strata (low-literacy: $(0.6,0.2,0.2)$; high-literacy: $(0.3,0.5,0.2)$). We estimate $\mathbb{E}_{\pi}[\beta_{w}R_{\text{tool}}+\gamma_{w}R_{\text{eng}}]$ with SNIPS and $\mathbb{E}_{\pi}[\alpha_{w}R_{\text{user}}]$ with AIPW using a rating-propensity correction. Because $R_{\text{obj}}$ (unweighted) and $R_{\text{total}}$ (weighted) use different compositions and estimators (SNIPS vs. AIPW), columns are reported on their native scales and do not sum algebraically in Table 5.
@@ 4.4 Counterfactual Policy Analysis
Table 5: Offline policy evaluation on deployment logs. $R_{\text{obj}}$: SNIPS (unweighted objective); $R_{\text{user}}$: AIPW (z-score). $R_{\text{total}}$: literacy-weighted combination. Brackets show session-level bootstrap 95% CIs for $R_{\text{total}}$.
Policy $R_{\text{obj}}$ (SNIPS) $R_{\text{user}}$ (AIPW) $R_{\text{total}}$